An egg is dropped from a third-floor window and lands on a foam rubber pad without breaking. If the egg has a mass of 49.3g and fell 10.7m from rest and the 6.25cm thick pad stops the egg in 7.48ms, how much does the pad get compressed
1
Expert's answer
2020-10-19T13:24:11-0400
"0.5mv^2=mgh\to v^2=2gh\\v^2=2(9.8)(10.7)\\v=14.5\frac{m}{s}\\mv=Ft\\(0.0493)(14.5)=(0.00748)F\\F=95.6\ N"